# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ 2. Pure spinor superparticle in 11D
where $\epsilon\to 0$ and the sign is - for commutators, and ++ for anticommutators. The action (2.2) is supplemented by the BRST charge $Q_{0}=\lambda^{\alpha}d_{\alpha}$, where $d_{\alpha}=p_{\alpha}-{1\over 2}(\gamma^{a}\theta)_{\alpha}P_{a}$ is the Brink-Schwarz fermionic constraint [29]. It has been shown that the quantization of this model reproduces all the fields and antifields of the Batalin-Vilkovisky description of linearized 11D supergravity [7]. In particular, the antifield of the ghost-for-ghost-for-ghost associated to the gauge symmetry in 11D supergravity, is located at the ghost number seven sector, and it represents the top cohomology of the BRST charge. Due to this, it has been suggested as a candidate for defining manifestly supersymmetric correlation functions in 11D, as we will discuss in next section.
@@ 2.2. Single-particle vertex operators
More recently, the ghost number zero vertex operator was constructed by using the physical operators studied above. It takes the strikingly compact form
@@ 2.3. Multi-particle vertex operators
$\eqalignno{&{{{\tenmsb U}}}^{(a)}=\sum_{P}U^{(a)}_{P}e^{ik_{P}\cdot X(z_{P})}=U^{(a)}_{i}e^{ik_{i}\cdot X(z_{i})}+U^{(a)}_{ij}e^{ik_{ij}\cdot X(z_{ij})}+\cdots,&({2.32})}$
where $a=0,1$ is the ghost number. We refer to the number of particles in a letter $P$ as the rank of the associated operator. After plugging the expansion (2.32) into eqns. (2.30) and (2.31), and properly matching the powers of the respective exponentials, one can find a set of relations obeyed by these operators for any specific value of the rank.
@@
$\eqalignno{&U_{ij}^{(1)}={1\over k_{ij}}\left[U^{(1)}_{i},U^{(0)}_{j}\right],&({2.34})}$
where $k_{ij}=k_{i}\cdot k_{j}$. Acting with the BRST charge and using the descent eqn. (2.26), it is not hard to see that (2.33) indeed holds. This solution can be shown to be unique up to BRST transformations. In addition, one can explicitly verify that the antisymmetric part of the operator (2.34) is BRST-trivial
$\eqalignno{&\left[Q,{1\over k^{2}_{ij}}\left[U^{(0)}_{i},U^{(0)}_{j}\right]\right]=\left(U^{(1)}_{ji}-U^{(1)}_{ij}\right).&({2.35})}$
@@
After plugging eqns. (2.25), (2.27) into the definition of the 2-rank operator (2.34), one finds that the pinching operator (2.36) can be expressed in the following compact way
@@
with $\tilde{\phi}_{23}^{a}=\phi_{23}^{a}-k_{23}\psi_{23}^{a}$, $\tilde{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-k_{23}\psi_{23}^{\alpha}$ and $\tilde{\phi}_{23,ab}=\phi_{23,ab}-k_{23}\psi_{23,ab}$, where $s_{ij}=(k_{i}+k_{j})^{2}=2k_{ij}$. These superfields correspond to two-particle configurations in 11D supergravity, and are explicitly constructed in Appendix B.2,
$\eqalignno{&\phi_{23}^{a}=\lambda^{\alpha}h_{23,\alpha}{}^{a},\ \ \phi_{23}^{\alpha}=\lambda^{\beta}h_{23,\beta}{}^{\alpha},\ \ (\lambda\gamma^{ab})^{\alpha}\phi_{23,ab}=\Omega_{23}^{\alpha},&({2.39})\cr&\psi_{23}^{a}=\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a},&({2.40})\cr&\psi_{23}^{\alpha}=\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha},&({2.41})\cr&(\lambda\gamma^{ab})^{\alpha}\psi_{23,ab}=-\phi^{a}_{(2}\Omega_{3),a}{}^{\alpha}+\phi^{c}_{(2}\lambda^{\beta}T_{3),\beta c}{}^{\alpha},&({2.42})}$
@@
$\eqalignno{&\{Q,\tilde{\phi}^{a}_{23}\}=\lambda\gamma^{a}\tilde{\phi}_{23}-k_{23}\big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\big{)},&({2.43})\cr&[Q,\tilde{\phi}^{\alpha}_{23}]=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}-k_{23}\big{(}\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\big{)}.&({2.44})}$
@@
$\eqalignno{&U_{ij}^{(3)}=\left[U^{(3)}_{i},U_{j}^{(0)}\right]-3k_{ij}\bigg{(}C_{i,a}\phi_{j}^{a}+C_{i,\alpha}\phi_{j}^{\alpha}\bigg{)}.&({2.45})}$
@@
$\eqalignno{&\ \ U_{ij}^{(3)}=3\Big{[}k_{i,a}\Omega_{i}^{ab}C_{j,b}-(\lambda\gamma_{ab}\lambda)h_{i}^{ab}k_{i,a}\phi_{j}^{c}+(\lambda\gamma_{a})_{\alpha}T_{i,b}{}^{\alpha}C_{j}^{ab}\Big{]}.&({2.46})}$
@@
$\eqalignno{&\{Q,U_{ij}^{(3)}\}=3k_{ij}(\lambda\gamma^{ab}\lambda)\phi^{a}_{i}\phi^{b}_{j}.&({2.47})}$
@@
$\eqalignno{&D_{\alpha}U^{(3)}_{12}+3\{Q,C_{\alpha,12}\}=-3(\gamma^{a}\lambda)_{\alpha}C_{a,12}&\cr&+3k_{12}\Big{(}2(\lambda\gamma_{ab})_{\alpha}\phi^{a}_{1}\phi_{2}^{b}+(\lambda\gamma_{ab}\lambda)h_{1,\alpha}^{a}\phi_{2}^{b}+(\lambda\gamma_{ab}\lambda)h_{2,\alpha}^{a}\phi_{1}^{b}\Big{)},&({2.48})\cr&\partial_{a}U^{(3)}_{12}-3[Q,C_{a,12}]=3(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}&\cr&+3k_{12}\Big{(}(\lambda\gamma_{ab}\phi_{1})\phi_{2}^{b}+(\lambda\gamma_{ab}\phi_{2})\phi_{1}^{b}+(\lambda\gamma_{bc}\lambda)\phi_{1}^{c}h^{b}_{2,a}+(\lambda\gamma_{bc}\lambda)\phi_{2}^{c}h^{b}_{1,a}\Big{)},&({2.49})}$
@@ 3.2. Simplification
The BRST transformations (3.8) induce a variation in the corresponding pinching operator $V_{ij}^{(1)}$ in the form dictated by eqn. (3.6). Since the correlator (3.1) has already been shown to be invariant under such transformations, one concludes that
$\eqalignno{&{\cal A}_{4}=\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle+\left\langle U_{1}^{(3)}(\infty)\tilde{V}_{23}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle,&({3.10})}$
@@
$\eqalignno{&\langle p_{m}(\tau_{1}){\cal O}(x(\tau_{2}))\rangle=\sigma_{12}\langle\partial_{m}{\cal O}(x(\tau_{2}))\rangle.&({3.11})}$
@@
$\eqalignno{&\int_{0}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle&\cr=&{1\over 2}\int_{0}^{\infty}d\tau\left\langle\Big{[}U_{1}^{(3)},\tilde{U}_{2}^{(0)}\Big{]}\tilde{U}_{3}^{(1)}U_{4}^{(3)}-U_{1}^{(3)}\Big{[}\tilde{U}_{2}^{(0)},\tilde{U}_{3}^{(1)}U_{4}^{(3)}\Big{]}\right\rangle\cr=&\int_{0}^{\infty}d\tau\left\langle\Big{[}U_{1}^{(3)},\tilde{U}_{2}^{(0)}\Big{]}\tilde{U}_{3}^{(1)}U_{4}^{(3)}\right\rangle,&({3.12})\cr}$
@@
$\eqalignno{&\int_{0}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle\cr&=\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\right],\tilde{U}_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.13})}$
@@
$\eqalignno{{\cal A}_{4}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\right],\tilde{U}_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle+\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}\left\{\tilde{U}_{3}^{(1)},\left[\tilde{U}_{2}^{(0)},U_{4}^{(3)}\right]\right\}\right\rangle\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left\langle\left\{U_{1}^{(3)},\tilde{V}_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.14})}$
@@
$\eqalignno{{\cal A}_{4}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},U_{2}^{(0)}\right],U_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle+\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}\left\{U_{3}^{(1)},\left[U_{2}^{(0)},U_{4}^{(3)}\right]\right\}\right\rangle\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left\langle\left\{U_{1}^{(3)},V_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.15})}$
@@
$\eqalignno{{\cal A}_{4}^{s}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},U_{2}^{(0)}\right],U_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle&\cr=&{6\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{b}_{2}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{1}^{c}h^{b}_{2,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{2}^{c}h^{b}_{1,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{2}^{a}h_{1,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{2,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle,&({3.16})}$
where $C_{ij}=C_{i,a}\phi_{j}^{a}-C_{i,\alpha}\phi_{j}^{\alpha}$, and $s_{ij}=(k_{i}+k_{j})^{2}=2k_{ij}$. A full derivation of (3.16) can be found in Appendix C.
@@
$\eqalignno{&\left\langle\left\{Q,\left(U_{1}^{(3)}\int_{-\infty}^{0}U_{3}^{(0)}\int_{0}^{\infty}U_{2}^{(0)}U_{4}^{(3)}\right)\right\}\right\rangle\cr=&-\left\langle U_{1}^{(3)}\int_{0}^{\infty}U_{2}^{(0)}U_{3}^{(1)}U_{4}^{(3)}\right\rangle+\left\langle U_{1}^{(3)}\int_{-\infty}^{0}U_{3}^{(0)}U_{2}^{(1)}U_{4}^{(3)}\right\rangle.&({3.17})}$
@@
$\eqalignno{{\cal A}_{4}^{t}=&\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}(\infty)\left\{U_{3}^{(1)}(0),\left[U_{2}^{(0)}(\tau),U_{4}^{(3)}(-\infty)\right]\right\}\right\rangle\cr=&\int_{0}^{\infty}d\tau\left\langle\Big{\{}\Big{[}U_{1}^{(3)}(\infty),U_{3}^{(0)}(\tau)\Big{]},U_{2}^{(1)}(0)\Big{\}}U_{4}^{(3)}(-\infty)\right\rangle\cr=&{6\over s_{13}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{13}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{13},U_{2}^{(1)}\}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{b}_{3}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{1}^{c}h^{b}_{3,a}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{3}^{c}h^{b}_{1,a}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}h_{1,\alpha}^{b}\phi_{2}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{3,\alpha}^{b}\phi_{2}^{\alpha}U_{4}^{(3)}\right\rangle.&({3.18})}$

### v2

@@ 2. Pure spinor superparticle in 11D
where $\epsilon\to 0$ and the sign is - for commutators, and ++ for anticommutators.
The action (2.2) is supplemented by the BRST charge $Q_{0}=\lambda^{\alpha}d_{\alpha}$, where $d_{\alpha}=p_{\alpha}-{1\over 2}(\gamma^{a}\theta)_{\alpha}P_{a}$ is the Brink-Schwarz fermionic constraint [29]. It has been shown that the quantization of this model reproduces all the fields and antifields of the Batalin-Vilkovisky description of linearized 11D supergravity [7]. In particular, the antifield of the ghost-for-ghost-for-ghost associated to the gauge symmetry in 11D supergravity, is located at the ghost number seven sector, and it represents the top cohomology of the BRST charge. Due to this, it has been suggested as a candidate for defining manifestly supersymmetric correlation functions in 11D, as we will discuss in next section.
@@ 2.2. Single-particle vertex operators
More recently, the ghost number zero vertex operator was constructed by using the physical operators studied above. It takes the strikingly compact form [27]
@@ 2.3. Multi-particle vertex operators
$\eqalignno{&{{{\tenmsb U}}}^{(a)}=\sum_{P}U^{(a)}_{P}e^{ik_{P}\cdot X(\tau)}=U^{(a)}_{i}e^{ik_{i}\cdot X(\tau)}+U^{(a)}_{ij}e^{ik_{ij}\cdot X(\tau)}+\cdots,&({2.32})}$
where $a=0,1$ is the ghost number. We refer to the number of particles in a letter $P$ as the rank of the associated operator. After plugging the expansion (2.32) into eqns. (2.30) and (2.31), and properly matching the powers of the respective exponentials, one can find a set of relations obeyed by these operators for any specific value of the rank. As usual, the action of the Hamiltonian on the exponential factors introduces $\tau$-derivatives in these relations.
@@
$\eqalignno{&U_{ij}^{(1)}={2\over s_{ij}}\left[U^{(1)}_{i},U^{(0)}_{j}\right],&({2.34})}$
where $s_{ij}=(k_{i}+k_{j})^{2}=2k_{i}\cdot k_{j}$. Acting with the BRST charge and using the descent eqn. (2.26), it is not hard to see that (2.33) indeed holds. This solution can be shown to be unique up to BRST transformations. In addition, one can explicitly verify that the antisymmetric part of the operator (2.34) is BRST-trivial
$\eqalignno{&\left[Q,{4\over s^{2}_{ij}}\left[U^{(0)}_{i},U^{(0)}_{j}\right]\right]=\left(U^{(1)}_{ji}-U^{(1)}_{ij}\right).&({2.35})}$
@@
After plugging eqns. (2.25), (2.27) into the definition of the rank-2 operator (2.34), one finds that the pinching operator (2.36) can be expressed in the following compact way
@@
with $\tilde{\phi}_{23}^{a}=\phi_{23}^{a}-{s_{23}\over 2}\psi_{23}^{a}$, $\tilde{\phi}_{23}^{\alpha}=\phi_{23}^{\alpha}-{s_{23}\over 2}\psi_{23}^{\alpha}$ and $\tilde{\phi}_{23,ab}=\phi_{23,ab}-{s_{23}\over 2}\psi_{23,ab}$. These superfields correspond to two-particle configurations in 11D supergravity, and are explicitly constructed in Appendix B.2,
$\eqalignno{&\phi_{23}^{a}=\lambda^{\alpha}h_{23,\alpha}{}^{a},\ \ \phi_{23}^{\alpha}=\lambda^{\beta}h_{23,\beta}{}^{\alpha},\ \ (\lambda\gamma^{ab})^{\alpha}\phi_{23,ab}=\Omega_{23}^{\alpha},&({2.39})\cr&\psi_{23}^{a}=\phi_{2}^{c}h_{3,c}{}^{a}+\phi_{3}^{c}h_{2,c}{}^{a}+\phi_{2}^{\delta}h_{3,\delta}{}^{a}+\phi_{3}^{\delta}h_{2,\delta}{}^{a},&({2.40})\cr&\psi_{23}^{\alpha}=\phi_{2}^{c}h_{3,c}{}^{\alpha}+\phi_{3}^{c}h_{2,c}{}^{\alpha}+\phi_{2}^{\delta}h_{3,\delta}{}^{\alpha}+\phi_{3}^{\delta}h_{2,\delta}{}^{\alpha},&({2.41})\cr&(\lambda\gamma^{ab})^{\alpha}\psi_{23,ab}=-2\phi^{a}_{(2}\Omega_{3),a}{}^{\alpha}+2\phi^{c}_{(2}\lambda^{\beta}T_{3),\beta c}{}^{\alpha},&({2.42})}$
@@
$\eqalignno{&\{Q,\tilde{\phi}^{a}_{23}\}=\lambda\gamma^{a}\tilde{\phi}_{23}-s_{23}\big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{a}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{a}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\big),&({2.43})\cr&[Q,\tilde{\phi}^{\alpha}_{23}]=(\lambda\gamma^{ab})^{\alpha}\tilde{\phi}_{23,ab}-s_{23}\big(\phi_{(2}^{c}\partial_{c}\phi_{3)}^{\alpha}+\phi_{(2}^{\delta}D_{\delta}\phi_{3)}^{\alpha}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\big).&({2.44})}$
@@
$\eqalignno{&U_{ij}^{(3)}=\left[U^{(3)}_{i},U_{j}^{(0)}\right]-{3s_{ij}\over 2}\bigg(C_{i,a}\phi_{j}^{a}+C_{i,\alpha}\phi_{j}^{\alpha}\bigg).&({2.45})}$
@@
$\eqalignno{&\ \ U_{ij}^{(3)}=3\Big[k_{i,a}\Omega_{i}^{ab}C_{j,b}-(\lambda\gamma_{cb}\lambda)h_{i}^{ab}k_{j,a}\phi_{j}^{c}+(\lambda\gamma_{a})_{\alpha}T_{i,b}{}^{\alpha}C_{j}^{ab}\Big].&({2.46})}$
@@
$\eqalignno{&\{Q,U_{ij}^{(3)}\}={3s_{ij}\over 2}(\lambda\gamma^{ab}\lambda)\phi^{a}_{i}\phi^{b}_{j}.&({2.47})}$
@@
$\eqalignno{&D_{\alpha}U^{(3)}_{12}+3\{Q,C_{\alpha,12}\}=-3(\gamma^{a}\lambda)_{\alpha}C_{a,12}&\cr&+{3s_{12}\over 2}\Big(2(\lambda\gamma_{ab})_{\alpha}\phi^{a}_{1}\phi_{2}^{b}+(\lambda\gamma_{ab}\lambda)h_{1,\alpha}^{a}\phi_{2}^{b}+(\lambda\gamma_{ab}\lambda)h_{2,\alpha}^{a}\phi_{1}^{b}\Big),&({2.48})\cr&\partial_{a}U^{(3)}_{12}-3[Q,C_{a,12}]=3(\lambda\gamma_{ba}\lambda)\phi^{b}_{12}&\cr&+{3s_{12}\over 2}\Big((\lambda\gamma_{ab}\phi_{1})\phi_{2}^{b}+(\lambda\gamma_{ab}\phi_{2})\phi_{1}^{b}+(\lambda\gamma_{bc}\lambda)\phi_{1}^{c}h^{b}_{2,a}+(\lambda\gamma_{bc}\lambda)\phi_{2}^{c}h^{b}_{1,a}\Big),&({2.49})}$
@@ 3.2. Simplification
Notice that (3.8) corresponds to gauge transformations of the form (3.6), where $\Lambda_{3}$ and $\Omega_{2}$ are determined by the BRST-exact terms enclosed in parentheses, while $\Lambda_{2}=0$. In this way, (3.8) generates a variation in the associated pinching operator $V_{ij}^{(1)}$ consistent with the structure prescribed by eqn. (3.6). The transformed operator then takes the explicit form
$\eqalignno{&\tilde{V}^{(1)}_{ij}=V_{ij}+[U^{(1)}_{i},\Lambda_{j}].&({3.10})}$
Since the correlator (3.1) has already been shown to be invariant under such transformations, one concludes that
$\eqalignno{&{\cal A}_{4}=\int_{-\infty}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle+\left\langle U_{1}^{(3)}(\infty)\tilde{V}_{23}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle,&({3.11})}$
@@
$\eqalignno{&\langle p_{m}(\tau_{1}){\cal O}(x(\tau_{2}))\rangle=\sigma_{12}\langle\partial_{m}{\cal O}(x(\tau_{2}))\rangle.&({3.12})}$
@@
$\eqalignno{&\int_{0}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle&\cr=&{1\over 2}\int_{0}^{\infty}d\tau\left\langle\Big[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\Big]\tilde{U}_{3}^{(1)}U_{4}^{(3)}-U_{1}^{(3)}\Big[\tilde{U}_{2}^{(0)},\tilde{U}_{3}^{(1)}U_{4}^{(3)}\Big]\right\rangle\cr=&\int_{0}^{\infty}d\tau\left\langle\Big[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\Big]\tilde{U}_{3}^{(1)}U_{4}^{(3)}\right\rangle,&({3.13})\cr}$
@@
$\eqalignno{&\int_{0}^{\infty}d\tau\left\langle U_{1}^{(3)}(\infty)\tilde{U}_{2}^{(0)}(\tau)\tilde{U}_{3}^{(1)}(0)U_{4}^{(3)}(-\infty)\right\rangle\cr&=\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\right],\tilde{U}_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.14})}$
@@
$\eqalignno{{\cal A}_{4}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},\tilde{U}_{2}^{(0)}\right],\tilde{U}_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle+\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}\left\{\tilde{U}_{3}^{(1)},\left[\tilde{U}_{2}^{(0)},U_{4}^{(3)}\right]\right\}\right\rangle\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left\langle\left\{U_{1}^{(3)},\tilde{V}_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.15})}$
@@
$\eqalignno{{\cal A}_{4}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},U_{2}^{(0)}\right],U_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle+\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}\left\{U_{3}^{(1)},\left[U_{2}^{(0)},U_{4}^{(3)}\right]\right\}\right\rangle\cr&\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left\langle\left\{U_{1}^{(3)},V_{23}^{(1)}\right\}U_{4}^{(3)}\right\rangle.&({3.16})}$
@@
$\eqalignno{{\cal A}_{4}^{s}=&\int_{0}^{\infty}d\tau\left\langle\left\{\left[U_{1}^{(3)},U_{2}^{(0)}\right],U_{3}^{(1)}\right\}U_{4}^{(3)}\right\rangle&\cr=&{6\over s_{12}}\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ba}\phi_{1})\phi^{b}_{2}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ba}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ba}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{1}^{c}h^{b}_{2,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{2}^{c}h^{b}_{1,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{2}^{a}h_{1,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{2,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle,&({3.17})}$
where $C_{ij}=C_{i,a}\phi_{j}^{a}-C_{i,\alpha}\phi_{j}^{\alpha}$. A full derivation of (3.17) can be found in Appendix C.
@@
$\eqalignno{&\left\langle\left\{Q,\left(U_{1}^{(3)}\int_{-\infty}^{0}U_{3}^{(0)}\int_{0}^{\infty}U_{2}^{(0)}U_{4}^{(3)}\right)\right\}\right\rangle\cr=&-\left\langle U_{1}^{(3)}\int_{0}^{\infty}U_{2}^{(0)}U_{3}^{(1)}U_{4}^{(3)}\right\rangle+\left\langle U_{1}^{(3)}\int_{-\infty}^{0}U_{3}^{(0)}U_{2}^{(1)}U_{4}^{(3)}\right\rangle.&({3.18})}$
@@
$\eqalignno{{\cal A}_{4}^{t}=&\int_{-\infty}^{0}d\tau\left\langle U_{1}^{(3)}(\infty)\left\{U_{3}^{(1)}(0),\left[U_{2}^{(0)}(\tau),U_{4}^{(3)}(-\infty)\right]\right\}\right\rangle\cr=&\int_{0}^{\infty}d\tau\left\langle\Big\{\Big[U_{1}^{(3)}(\infty),U_{3}^{(0)}(\tau)\Big],U_{2}^{(1)}(0)\Big\}U_{4}^{(3)}(-\infty)\right\rangle\cr=&{6\over s_{13}}\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{13}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{13},U_{2}^{(1)}\}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ba}\phi_{1})\phi^{b}_{3}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ba}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ba}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{1}^{c}h^{b}_{3,a}\phi_{2}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{3}^{c}h^{b}_{1,a}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr+&3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}h_{1,\alpha}^{b}\phi_{2}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{3,\alpha}^{b}\phi_{2}^{\alpha}U_{4}^{(3)}\right\rangle.&({3.19})}$